OpenStudy (anonymous):

Simplify. square root of 9 over sixth root of 9 9 to the power of negative 2 over 3 9 to the power of negative 1 over 3 8 to the power of 3 over 10 9 to the power of 1 over 3

OpenStudy (anonymous):

\[\Large \frac{ \sqrt{9} }{ \sqrt[6]{9} }\]correct?

OpenStudy (zzr0ck3r):

\(\frac{a^b}{a^c}=a^{b-c}=\frac{1}{a^{c-b}}\\\sqrt[a]{b^c}=b^{\frac{c}{a}}\\(\frac{a}{b})^{-c}=(\frac{b}{a})^c\)

OpenStudy (jdoe0001):

keep in mind that \(\Large \bf \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}= \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}} \)

OpenStudy (zzr0ck3r):

so for the first one \(\frac{\sqrt{9}}{\sqrt[6]{9}}=\frac{9^{\frac{1}{2}}}{9^{\frac{1}{6}}}=9^{\frac{1}{2}-\frac{1}{6}}=9^{\frac{2}{6}}=\sqrt[6]{9^2}\)

OpenStudy (zzr0ck3r):

there must be an echo in here

OpenStudy (zzr0ck3r):

:)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Okay, thank you :)